1. (25 points) The market for study desks is characterized by perfect competition. Firms and consumers are price takers and in the long run there is free entry and exit of firms in this industry. All firms are identical in terms of their technological capabilities. Thus the cost function as given below for a representative firm can be assumed to function faced by each firm in the industry. The total cost and marginal cost functions t the representative firm are given by the following equations: The market demand is given by Po 1025-20 Note: Q represents market quantity and q represents an individual firm's output level.
(c) When this industry is in long-run equilibrium, how many firms are in the industry? (firms are identically sized). (d) Now suppose that the number of students increases such that the market demand curve for study desks shifts out and is given by Po-1525-20
Answer #1

(c) In long run equilibrium, P = MC = ATC, from which we obtain

q = 5

P = 25

From market demand function,

1,025 - 2Q = 25

2Q = 1,000

Q = 500 (market output)

Number of firms = Q / q = 500 / 5 = 100

NOTE: Part (d) question is incomplete.